# Robust Bayesian Scene Reconstruction by Leveraging Retrieval-Augmented Priors

### v1

@@ Abstract
Constructing 3D representations of object geometry is critical for many downstream manipulation tasks. These representations must be built from potentially noisy partial observations. In this work we focus on the problem of reconstructing a multi-object scene from a single RGBD image. Current deep learning approaches to this problem can be brittle to noisy real world observations and out-of-distribution objects. Other approaches that do not rely on training data cannot accurately infer the backside of objects. We propose BRRP, a reconstruction method that can leverage preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction. In order to make our method more efficient, we introduce the concept of retrieval-augmented prior, where we retrieve relevant components of our prior distribution during inference. Our method produces a distribution over object shape that can be used for reconstruction or measuring uncertainty. We evaluate our method in both procedurally generated scenes and in real world scenes. We show our method is more robust than a deep learning approach while being more accurate than a method with an uninformative prior.
@@ I Introduction
The ability to construct internal representations of its operating environment is key for robot autonomy. These representations need to be particularly fine-grained for robotic manipulation, which often requires closely interacting with and avoiding objects. These interactions make it necessary for robots to develop an understanding of the geometry within their vicinity. Explicit 3D representations of the geometry of the scene are often required for the robust usage of downstream grasping and motion planning algorithms. These representations must be built from observations that are both noisy and only contain partial information of the scene. In our case, we focus on the problem of building a 3D representation of multi-object scenes from a single RGBD camera image.
One approach to this problem is to build some type of neural network that predicts the full geometry of an object given a partial view. Such approaches are able to use existing mesh datasets to more accurately infer the backside of objects. Unfortunately, these approaches also tend to have a number of problems when used on the real world depth cameras. The presence of unknown objects, significant occlusion, noisy point clouds, or inaccurate segmentation prevent these methods from being reliably deployed in less-structured environments.
Most deep learning methods for 3D reconstruction also often lack a sense of uncertainty about the shape of the object. Such uncertainty has been shown to be useful in the downstream task of robotic grasping [1, 2]. Such uncertainty can also be useful in developing safe motion plans or during active learning.
Another common approach to building full 3D representations from a partial view is to simply build such a representation in a way that does not rely on prior information from mesh datasets. A common example of this is the Gaussian Process Implicit Surface [3] model. A more recent example is V-PRISM [4], which probabilistically maps tabletop scenes without using prior information. These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution. However, this also means that they don’t always accurately reconstruct the backside of objects.
@@
In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors. Our method is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects. In order to improve efficiency, we introduce the idea of a retrieval-augmented prior, inspired by [5], where we retrieve relevant components of our prior distribution based on classification results. We refer to our method as BRRP (pronounced ’burp’), which stands for Bayesian Reconstruction with Retrieval-augmented Priors. BRRP starts with an observed RGBD image with corresponding instance segmentations. Then, we compute an identification result that predicts which objects from our database should be retrieved to use as a prior during reconstruction. We then negatively sample the observation and use Stein Variational Gradient Descent (SVGD) [6] to infer a posterior distribution over object shapes. This distribution is defined as a set of Hilbert Maps [7] for each object. Because we solve for a distribution over weights, we can recover principled uncertainty about the shape of each object. In practice we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant. An example scene and reconstruction can be seen in Figure 1.
We conduct experiments on BRRP in both procedurally generated scenes and in the real world. We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects on the generated scenes. We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera. We also qualitatively examine the uncertainty from our method’s reconstructions.
@@ II Related Works
Explicit 3D Representations. There are many different ways of representing 3D geometry of a scene. In the mapping literature, techniques such as truncated signed distance functions [8] build voxelized representations of an environment. Hilbert maps [7], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space. Hilbert maps representations have also been extended to Bayesian Hilbert maps of various forms [9, 10, 11]. Neural implicit functions have also been used to represent continuous 3D geometry [12, 13, 14]. Other representations are built using differentiable rendering and combining multiple views. Neural radiance fields [15] are an example of this, in which a neural network maps 3D position to density and color. 3D Gaussian splatting [16] does a similar thing but with a set of Gaussians instead of a neural network. Another representation that has been studied is super quadrics [17].
@@
3D Reconstruction with Deep Learning. Many methods have been proposed as ways to leverage deep learning to reconstruct scenes or objects. While some methods aim to predict object shape from RGB data only [18, 19, 20, 21], we instead focus on using depth measurements during reconstruction. DeepSDF [12] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function. In the context of robotics, [22] extends DeepSDF to have uncertainty-awareness. Other work, such as occupancy networks [13] or PointSDF [23] try to directly predict such a latent code without inference-time optimization. Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [24, 25]. Language is also used during reconstruction in [26] and [27]. Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [28]. While these work typically focus on single object scenes, other work focuses on reconstruction scenes with multiple, occluding objects. For example, [29] learns a reconstruction from a voxel representation with different channels to account for occlusion. Another method uses silhouettes to refine the initial reconstructions [30]. In practice these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras. Inaccurate segmentation can also be a problem for many of these methods as well.
@@
3D Reconstruction without Deep Learning. There are also many approaches to perform probabilistic 3D reconstruction without deep learning. Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [31]. Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [3]. While there is an extension of GPIS to a slightly more informative prior [32], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors. In our work, we derive our prior from pre-existing mesh datasets. V-PRISM [4] is another method that probabilistically maps the scene using a multiclass framing.
@@ III-A Hilbert Maps
In order to recover the weight vector corresponding to a given depth observation, negative sampling is performed along the unoccupied portions of the depth rays. These negative samples are assigned a label of unoccupied and the points at the end of the ray are labeled as occupied. Then, stochastic gradient descent (SGD) is performed on the binary cross entropy of the negative samples and terminal points of the ray. The binary cross entropy measures the likelihood of the samples, and is defined as:
@@ III-B Stein Variational Gradient Descent
Stein Variational Gradient Descent (SVGD) [6] is a algorithm for variational inference that closely resembles gradient descent. The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$. Usually, this takes the form of an optimization problem over the Kullback-Leibler (KL) divergence:
@@
The particles can then be iteratively updated according to Equation 2:
@@ IV Retrieval-Augmented Priors
Retrieval-augmented generation [5] was originally introduced in the context of improving language generation. The work has served as inspiration for an approach to affordance-prediction in [36]. In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior distributions during Bayesian inference.
@@
$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$
@@ V The BRRP Method
Our method takes a single RGBD image and produces reconstructions for each object in the scene. We treat the problem as a Bayesian inference problem over an observation described by negative samples. We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV. We use CLIP to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh. We then use SVGD to optimize for a set of samples over map weights. We can generate predicted reconstructions by taking the expected occupancy over our weights for a given location. Figure 3 shows a visual overview of our method.
@@ V-B Retrieval-Augmented Priors for Hilbert Maps
Because it would be inefficient to register all meshes that are part of the prior mixture model, we propose using the retrieval-augmented prior approach introduced in Section IV. In order to determine which objects to use, we need to compute $P(c|R)$. In our case, we use CLIP [39] as a zero-shot classifier for our different objects. For each object in our prior, we store a small textual description of the object. These descriptions are then used as classes for CLIP to classify each segmented object. In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of the object and feed the cropped image for each object as the input into CLIP.
@@ VI-A Experimental Details
Baselines: We compare our work against two main baselines, V-PRISM [4] and a version of PointSDF [23] that predicts occupancy and is trained on ShapeNet [42] scenes. V-PRISM is a probabilistic mapping method that uses an uninformative prior. This means that it is robust to novel objects, but doesn’t accurately reconstruct object backsides. We refer to this baseline as V-PRISM. In contrast, PointSDF is a learning based method. This means it can leverage prior information from pre-existing mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift. We refer to this baseline as PointSDF. When computing reconstruction meshes with PointSDF, a level set of $\tau=0.3$ is used.
@@
Method ShapeNet Scenes YCB Scenes Objaverse Scenes
V-PRISM [4] 0.3092 0.5003 0.4640
PointSDF [23] 0.3600 0.4601 0.3471
BRRP (ours) 0.3124 0.5277 0.4809
@@ VI-B Results
Insight 1: BRRP is more accurate than a method with a uninformative prior.
As showcased in Table I, BRRP outperforms V-PRISM on each set of procedurally generated scenes. It has the highest IoU improvement when compared against V-PRISM on the YCB scenes. A similar pattern can be seen in the chamfer distances in Figure 5, where BRRP consistently outperforms V-PRISM, with the biggest gap occuring on the YCB scenes. This makes sense because BRRP uses a subset of the YCB objects as its prior.
We can see this improvement qualitatively in Figure 6. While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a portion of space that the object doesn’t occupy. BRRP is able to more accurately reconstruct these objects. The clearest example is the dark green object in the right-most scene.
@@
While PointSDF outperformed BRRP on the ShapeNet scenes on IoU, BRRP had a lower Chamfer distance. BRRP also performed better on mesh datasets that PointSDF was not trained on. In Table I, we can see that on Objaverse scenes, where the objects were novel to both methods, BRRP performed better than PointSDF. When measuring chamfer distance, BRRP outperformed PointSDF on all datasets as shown in Figure 5. These results suggests BRRP is more robust to different object distributions than PointSDF.
@@
On the real world scenes in Figure 6, BRRP is qualitatively more robust than PointSDF. PointSDF struggles with the noise associated with real world scenes as well as the novel objects. It tends to predict a spherical object on many non-spherical objects in the real world scenes. Our method on the other hand, is better able to reconstruct the objects on these scenes, including ones that are out-of-distribution for its prior.
@@
Figure 8: Visualization of surface uncertainty from BRRP. Lighter areas correspond to higher uncertainty about the shape.
@@ VII Conclusion
We introduced the concept of retrieval-augmented priors, where we retrieve relevant components of a prior distribution during Bayesian inference. We also introduced a novel Bayesian method for scene reconstruction that uses informative priors. Our method, BRRP, leveraged existing mesh datasets to build its prior and probabilistically reconstructed the scene leveraging these priors in a retrieval-augmented manner. We experimentally showed our method was more robust than a deep learning method as well as more accurate than a mapping method. Finally, showed a qualitative example of recovering principled uncertainty from our method.

### v2

@@ Abstract
Constructing 3D representations of object geometry is critical for many downstream robotics tasks, particularly tabletop manipulation problems. These representations must be built from potentially noisy partial observations. In this work, we focus on the problem of reconstructing a multi-object scene from a single RGBD image, generally from a fixed camera in the scene. Traditional scene representation methods generally cannot infer the geometry of unobserved regions of the objects from the image. Attempts have been made to leverage deep learning to train on a dataset of observed objects and representations, and then generalize to new observations. However, this can be brittle to noisy real-world observations and objects not contained in the dataset, and cannot reason about their confidence. We propose BRRP, a reconstruction method that leverages preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction. In order to make our method more efficient, we introduce the concept of retrieval-augmented prior, where we retrieve relevant components of our prior distribution during inference. The prior is used to estimate the geometry of occluded portions of the in-scene objects. Our method produces a distribution over object shape that can be used for reconstruction or measuring uncertainty. We evaluate our method in both simulated scenes and in the real world. We demonstrate the robustness of our method against deep learning-only approaches while being more accurate than a method without an informative prior.
@@ I Introduction
The ability to construct internal representations of its operating environment is key for robot autonomy. These representations need to be particularly fine-grained for robotic manipulation, which often requires closely interacting with and avoiding objects. These interactions make it necessary for robots to develop an understanding of the geometry within their vicinity. Explicit 3D representations of the geometry of the scene are often required for the robust usage of downstream grasping and motion planning algorithms. These representations must be built from observations that are both noisy and, due to occlusion, only contain partial information of the scene. In our case, we focus on the problem of building a 3D representation of multi-object scenes from a single RGBD camera image.
One approach to this problem is to train a neural network to predict the full geometry of an object given a partial view. Such approaches are able to use existing mesh datasets to more accurately infer the occluded backside of objects. Unfortunately, these approaches also tend to have a number of problems when used on real-world depth cameras. The presence of unknown objects, significant occlusion, noisy point clouds, or inaccurate segmentation prevents these methods from being reliably deployed in less-structured environments.
Deep learning methods for 3D reconstruction generally lack a sense of uncertainty about the shape of the object. This can be particularly detrimental when an object in the scene is only partially observed and we seek to factor in the geometry of observed and occluded regions. Such uncertainty can enable safer and more robust operation in a range of downstream tasks, such as robot grasping [1, 2], safe motion generation and active learning.
Another common approach to building full 3D representations from a partial view is to build a representation, solely from the observed data without considering prior information from mesh datasets. A common example of this is the Gaussian Process Implicit Surface [3] model. A more recent example is V-PRISM [4], which probabilistically maps tabletop scenes without using prior information. These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution. However, this also means that they cannot reconstruct the unobserved backside of objects. Humans have the remarkable ability to infer the geometry of scenes based on prior experience, we seek to imbue robots with the same capability.
@@
In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors. We present Bayesian Reconstruction with Retrieval-augmented Priors (BRRP). BRRP is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects. To further improve efficiency, we introduce the idea of a retrieval-augmented prior, where we retrieve relevant components of our prior distribution based on classification results. We begin with an observed RGBD image with corresponding instance segmentations. Then, we compute an identification result that predicts which objects from our database should be retrieved to use as a prior during reconstruction. We use this prior along with a sampled likelihood to infer a posterior distribution over object shapes. Because we solve for a distribution, we can recover principled uncertainty about each object’s shape. In practice, we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant. An example scene and reconstruction can be seen in Figure 1.
We conduct experiments on BRRP in both procedurally generated scenes and in the real world. We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects in the generated scenes. We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera, and is able to capture uncertainty within the reconstructions.
@@ II Related Works
3D Representations. There are many different ways of representing 3D geometry of a scene. In the mapping literature, techniques such as truncated signed distance functions [5] build voxelized representations of an environment. Hilbert maps [6], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space. Hilbert map representations have also been extended to Bayesian Hilbert maps of various forms [7, 8, 9]. Neural implicit functions have also been used to represent continuous 3D geometry [10, 11, 12]. Other representations are built using differentiable rendering and combining multiple views, especially for novel-view synthesis. A Neural radiance field (NeRF) [13] is an example of this, in which a neural network maps 3D position to density and color. 3D Gaussian splatting [14] tackles a similar but with a set of Gaussians instead of a neural network. Foundation models have also been developed for this task in [15] and applied to robotics [16]. Other representation primitives have been studied, including super quadrics [17].
@@
3D Reconstruction with Deep Learning. Many methods have been proposed as ways to leverage deep learning to reconstruct scenes or objects. While some methods aim to predict object shape from RGB data only [18, 19, 20, 21], we instead focus on using depth measurements during reconstruction. DeepSDF [10] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function. In the context of robotics, [22] extends DeepSDF to have uncertainty-awareness. Other work, such as occupancy networks [11] or PointSDF [23] try to directly predict such a latent code without inference-time optimization. Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [24, 25]. Language is also used during reconstruction in [26] and [27]. Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [28]. While these works typically focus on single object scenes, other work focuses on reconstruction scenes with multiple occluding objects. For example, [29] learns a reconstruction from a voxel representation with different channels to account for occlusion. Another method uses silhouettes to refine the initial reconstructions [30]. In practice, these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras. Inaccurate segmentation can also be a problem for many of these methods as well.
@@
3D Reconstruction without Deep Learning. There are also many approaches to perform probabilistic 3D reconstruction without deep learning. Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [31]. Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [3]. While there is an extension of GPIS to a slightly more informative prior [32], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors. In our work, we derive our prior from pre-existing mesh datasets. V-PRISM [4] is another method that probabilistically maps the scene using a multi-class framing.
@@ III-A Hilbert Maps
To recover the weight vector corresponding to a given depth observation, negative sampling is performed along the unoccupied portions of the depth rays. These negative samples are assigned a label of unoccupied and the points at the end of the ray are labeled as occupied. Then, stochastic gradient descent (SGD) is performed on the binary cross entropy (BCE) of the negative samples and terminal points of the ray. The binary cross entropy measures the likelihood of the samples, and is defined as:
@@ III-B Stein Variational Gradient Descent
Stein Variational Gradient Descent (SVGD) [36] is an algorithm for variational inference that closely resembles gradient descent. The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$. Usually, this takes the form of an optimization problem over the Kullback-Leibler (KL) divergence:
@@
The particles can then be iteratively updated according to $\hat{g}^{*}$ in Equation 2 with:
@@ IV Retrieval-Augmented Priors
Retrieval-augmented generation [37] was originally introduced in the context of improving language generation. The work has served as inspiration for an approach to affordance-prediction in [38]. In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior density functions during Bayesian inference.
@@
$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$ (4)
@@ V The BRRP Method
Our method takes a single RGBD image and produces reconstructions for each object in the scene. We treat the problem as a Bayesian inference problem over an observation described by negative samples. We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV. We use CLIP [39] to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh. We then use SVGD to optimize for a set of samples over map weights. We can generate predicted reconstructions by taking the expected occupancy over our weights for a given location. Figure 3 shows a visual overview of our method.
@@ V-B Retrieval-Augmented Priors for Hilbert Maps
Because it would be inefficient to register all meshes that are part of the prior mixture model, we propose using the retrieval-augmented prior approach introduced in Section IV. In order to determine which objects to use, we need to compute $P(c|R)$ from Equation 4. In our case, we use CLIP [39] as a zero-shot classifier for our different objects. For each object in our prior, we store a small textual description of the object. These descriptions are then used as classes for CLIP to classify each segmented object. In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of each object and feed the cropped images as input into CLIP.
@@ VI-A Experimental Details
Baselines: We compare our work against two main baselines, V-PRISM [4] and a version of PointSDF [23] that predicts occupancy and is trained on ShapeNet [44] scenes. V-PRISM is a probabilistic mapping method that uses an uninformative prior. This means that it is robust to novel objects, but doesn’t accurately reconstruct object backsides. We refer to this baseline as V-PRISM. In contrast, PointSDF is a learning-based method. This means it can leverage prior information from mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift. We refer to this baseline as PointSDF. When computing reconstruction meshes with PointSDF, a level set of $\tau=0.3$ is used.
@@
Method ShapeNet Scenes YCB Scenes Objaverse Scenes V-PRISM [4] 0.3092 0.5003 0.4640 PointSDF [23] 0.3600 0.4601 0.3471 BRRP (ours) 0.3124 0.5277 0.4809
@@ VI-B Results
Insight 1: BRRP is more accurate than a method with an uninformative prior.
As showcased in Table I, BRRP outperforms V-PRISM on each set of procedurally generated scenes. It has the highest IoU improvement from V-PRISM on the YCB scenes. A similar pattern can be seen in the chamfer distances in Figure 5, where BRRP consistently outperforms V-PRISM, with the biggest gap of 0.19 (cm) occuring on the YCB scenes. This makes sense because BRRP uses a subset of the YCB objects as its prior.
We can see this improvement qualitatively in Figure 6. While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a large portion of space that the object doesn’t occupy. BRRP is able to more accurately reconstruct these objects. The clearest example of this is the dark green object in the right-most scene in Figure 6.
@@
While PointSDF outperformed BRRP on the ShapeNet scenes on IoU (Table I), BRRP had a lower Chamfer distance (Figure 5). BRRP also performed better on mesh datasets that PointSDF was not trained on. In Table I, we can see that on Objaverse scenes, where the objects were novel to both methods, BRRP performed better than PointSDF. When measuring chamfer distance, BRRP outperformed PointSDF on all datasets as shown in Figure 5. These results suggests BRRP is more robust to different object distributions than PointSDF.
@@
On the real world scenes in Figure 6, BRRP is qualitatively more robust than PointSDF. PointSDF struggles with the noise associated with real world scenes as well as the novel objects. It tends to predict a spherical object on many non-spherical objects in the real world scenes. Our method on the other hand, is better able to reconstruct these scenes, including the objects that are out-of-distribution for its prior.
@@
Figure 8: Visualization of a cylindrical Clorox container surface uncertainty from BRRP. Lighter areas correspond to higher uncertainty about the shape. We observe that the occluded back-side of the container has high uncertainty.
@@ VII Conclusion
We introduced the concept of retrieval-augmented priors, where we retrieve relevant components of a prior distribution during Bayesian inference. We also introduced a novel Bayesian method for scene reconstruction that uses informative priors. Our method, BRRP, leveraged existing mesh datasets to build its prior and probabilistically reconstructed the scene leveraging these priors in a retrieval-augmented manner. We experimentally showed our method was more robust than a deep learning method as well as more accurate than a mapping method. Finally, we showed a qualitative example of recovering principled uncertainty from our method.